OpenStudy (anonymous):

Solve the system of equations by graphing 3x + 2y = -2 -4x + 5y = -28

OpenStudy (anonymous):

Do you have a graphing calculator or other utilities?

OpenStudy (anonymous):

Or do they want you to actually draw it. (Either way should be fine)

OpenStudy (geometry_hater):

You need to graph it and then look for a point of its intersection

OpenStudy (anonymous):

Unfortunately I do not have a graphing calculator.

OpenStudy (geometry_hater):

You can download one onto your computer

OpenStudy (anonymous):

I would just use wolframalpha. The way these equations work out, there's no way you'd get an accurate answer on pencil/paper.

OpenStudy (geometry_hater):

or use a free one here http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html

OpenStudy (geometry_hater):

So to start you need to get y by itself in both equations

OpenStudy (anonymous):

If you were using a TI calculator or some similar, you would have to isolate y.

OpenStudy (anonymous):

How do I isolate y?

OpenStudy (anonymous):

Get y by itself in both equations.

OpenStudy (geometry_hater):

3x + 2y = -2 Subtract 3x ......2y = -3x -2 Divide by 2........ y= -3/2x - 1 Do the same for the other one

OpenStudy (anonymous):

Oh okay. Let me work that out.